The demand for football tickets is Q = 360 − 10P and the supply of football tickets is Q = 20P. Calculate the gross price paid by consumers after a per-ticket tax of $2. Calculate the after-tax price received by ticket sellers.
Answer #1

Answer

Converting the supply curve to inverse supply curve to add the tax

Q=20P

20P=Q

P=0.05Q

adding tax

P=0.05Q+2

converting back

P=0.05Q+2

0.05Q=P-2

Q=20P-40

the equilibrium is at

after-tax supply equal to the demand

20P-40=360-10P

30P=400

P=13.3333333=13.33=the gross price paid by consumers after a per-ticket tax

the after-tax price received by ticket sellers=P-per unit tax =13.33-2=11.33